OpenStudy (anonymous):

Solve equation 0.6 (y+3) = 4.8

OpenStudy (ttop0816):

0.6(y + 3) = 4.8 Reorder the terms: 0.6(3 + y) = 4.8 (3 * 0.6 + y * 0.6) = 4.8 (1.8 + 0.6y) = 4.8 Solving 1.8 + 0.6y = 4.8 Solving for variable 'y'. Move all terms containing y to the left, all other terms to the right. Add '-1.8' to each side of the equation. 1.8 + -1.8 + 0.6y = 4.8 + -1.8 Combine like terms: 1.8 + -1.8 = 0.0 0.0 + 0.6y = 4.8 + -1.8 0.6y = 4.8 + -1.8 Combine like terms: 4.8 + -1.8 = 3 0.6y = 3 Divide each side by '0.6'. y = 5 Simplifying y = 5